# Beyond Synthetic Benchmarks: Evaluating LLM Performance on Real-World Class-Level Code Generation

### v1

@@ 3.1. Dataset Curation
Table 1. An example Python class and corresponding class skeleton from codenerix/django-codenerix (githubGitHubTkaemmingdjangosubdomains, ).
@@
Following ClassEval’s exclusion criteria (du2023classeval, ), we omit tasks with complex execution dependencies, identified through static metrics. However, unlike ClassEval, we do not filter by domain, retaining a diverse range of topics such as Game Development, File Handling, and Management Systems. This approach ensures our dataset reflects the heterogeneity of real-world projects, providing a robust foundation for evaluating LLM performance on class-level code generation. The complete data curation process is shown in Figure 1.
Start Mine pre-cutoff classes from CodeSearchNet projects Mine post-cutoff classes from repositories created after March 31, 2025 Merge datasets (‘pre’ and ‘post’ falgs added for easy tracking) Filter engineered projects Static analysis with Understand™ Extract class skeletons using ast library Remove problematic classes Randomly sample 200 classes per version RealClassEval Dataset
@@ 3.2.1. Selected LLMs
To evaluate the performance of LLMs on class-level code generation, we selected seven state-of-the-art models: Qwen 2.5 Coder (hui2024qwen2, ), GPT-4.1 (openaiIntroducingGPT41, ), GPT-5 (openaiIntroducingGPT5, ), GPT-OSS (agarwal2025gpt, ), Codestral (mistralCodestralMistral, ), DeepSeek-V3 (liu2024deepseek, ), and Llama-4 Maverick (metaLlamaHerd, ). These models were chosen to represent a diverse range of architectures, training paradigms, and capabilities, ensuring a comprehensive evaluation of both general-purpose and code-specific LLMs, as well as models with varying reasoning abilities. They also vary in their release dates, knowledge cutoffs, parameter sizes, context window lengths, and specialization. Table 2 summarizes these characteristics, highlighting the diversity of our evaluation set.
@@
Qwen 2.5 Coder September 2024 June 2024 72B 128K Code-specific, optimized for code generation and reasoning
@@
Codestral May 2024 Unknown 22B 256K Code-specific, specialized for code, lightweight and efficient
@@ 3.2.2. Prompt Design
Unlike ClassEval (du2023classeval, ), which explores three generation strategies (holistic, incremental, and interactive), we focus exclusively on holistic generation. This decision is motivated by the findings of ClassEval, which indicate that models with longer context windows (such as GPT-3.5 and GPT-4 in their study) perform well in holistic generation. Given that all our selected models have substantial context windows (ranging from 128K to 1M tokens), we opt for holistic generation to leverage their ability to process and generate entire class implementations in a single pass.
@@ 3.3. Test Suite Generation
To assess functional correctness, we use PYNGUIN (lukasczyk2022pynguin, ) to automatically generate test suites for the classes in our dataset. Given the diverse nature of real-world projects, we lack sufficient domain knowledge to manually create meaningful test cases, limiting our ability to measure correctness beyond structural similarity. PYNGUIN, with its DynaMOSA (panichella2017automated, ) algorithm, is a suitable solution as it automatically generates test cases with high branch coverage, enabling us to evaluate the functional correctness of LLM-generated classes without manual intervention.
@@ 3.4.1. Performance Analysis
To evaluate the performance of LLMs on class-level code generation, we employ pass rate as our primary metric, defined as the ratio of passed tests to the total number of tests for each generated class. This metric is directly related to the widely used Pass@1 (chen2021evaluating, ), which measures the probability that a generated solution passes all tests in a single attempt. While Pass@1 is typically used for function-level evaluation, our pass rate extends this concept to class-level correctness, providing a robust measure of functional accuracy across the entire class implementation. For a given class $i$, the pass rate metric is defined as:
@@ 4.1.2. Kruskal-Wallis Test
We used the Kruskal-Wallis test (kruskal1952use, ) to compare per-class pass rates across the three datasets (ClassEval, Pre-Cutoff, Post-Cutoff) for each LLM. In the current context, the null hypothesis states that the pass rate distributions are identical across datasets, while the alternative hypothesis states that at least one dataset’s distribution differs.
@@
The Kruskal-Wallis test is a non-parametric alternative to one-way ANOVA, suitable for non-normal data and independent samples across three or more groups. The non-normality of pass rates was confirmed by the Shapiro-Wilk test (shapiro1965analysis, ), justifying the non-parametric approach.
@@ 4.1.3. Mann-Whitney U Tests
For significant Kruskal-Wallis results, we conducted pairwise Mann-Whitney U tests (mann1947test, ) (ClassEval vs. Pre-Cutoff, ClassEval vs. Post-Cutoff, Pre-Cutoff vs. Post-Cutoff) to identify specific differences. The Mann-Whitney U test is a non-parametric approach suitable for unpaired samples (distinct classes across datasets) and non-normal pass rates. We applied Benjamini-Hochberg (benjamini1995controlling, ) False Discovery Rate (FDR) correction globally across all 21 tests (3 comparisons per LLM, 7 LLMs) to control for multiple comparisons. The pairwise tests pinpoint which datasets differ, while FDR correction balances Type I error across all tests.
@@ 4.1.4. Effect Sizes and Confidence Intervals
We computed Cliff’s Delta (cliff1993dominance, ; macbeth2011cliff, ) to quantify the magnitude of pairwise differences, with thresholds (cohen2013statistical, ; rahman2024automatic, ): negligible (¡0.147), small (¡0.33), medium (¡0.474), large ($\geq$0.474). Bootstrap 95% confidence intervals (CIs) were calculated for Cliff’s Delta to quantify uncertainty, using 1,000 iterations. Cliff’s Delta provides an effect size for ordinal data, and CIs ensure robust estimation of differences. Although Cliff’s Delta help us get the effect size of differences, it does not quantify the actual differences. Hence, we also calculate the mean differences between the distributions with bootstrap 95% CIs for an easy interpretation.
@@ 4.1.5. Pooled Analysis
We performed a chi-square analysis (pearson1900x, ; fisher1922interpretation, ) on pooled test case counts (passed vs. failed) using 2×2 contingency tables for pairwise dataset comparisons. The chi-square test is appropriate for categorical outcomes. Cramér’s V (cramer1999mathematical, ) quantified effect sizes (cohen2013statistical, ): negligible (¡0.1), small (¡0.3), medium (¡0.5), large ($\geq$0.5), with bootstrap CIs for proportion differences. FDR correction was applied globally across all 21 tests. The pooled analysis provides an aggregate view of performance, complementing per-class results and ensuring robustness.
@@ 4.3. Interpretation
The high pass rate of ClassEval compared to RealClassEval was significant, but the reasons were initially unclear, as the large effect sizes (Cliff’s Delta 0.46–0.64) and pass rate differences did not fully explain the gap. To investigate, we conducted a mixed-effects model (laird1982random, ; carey2001mixed, ) analysis, including dataset, test count, and LLM-generated code metrics (lines of code, method count, cyclomatic complexity, and class coupling) as predictors, with LLM as a random effect. We chose these $4$ static metrics because they represent both the length of the code as well as the complexity of the code. The model showed significant dataset effects (e.g., $\beta=-1.060$, $p<0.001$) and a negative test count effect ($\beta=-0.014$, $p<0.001$). While lines of code ($\beta=-0.002$, $p=0.006$) and cyclomatic complexity ($\beta=0.043$, $p<0.001$) showed statistical significance, their effects were small relative to the dataset effect, indicating that code complexity metrics explain only a minor portion of the performance gap.
@@ 5.1.2. Friedman Test
Given the continuous and bounded nature of our outcome variable (pass rate $\in[0,1]$) and the paired structure of our data, we employed non-parametric statistical methods. The Friedman test (friedman1937use, ; friedman1940comparison, ) compared pass rates across three docstring conditions (full, partial, no) for each model-dataset combination, testing the null hypothesis that pass rate distributions are identical versus the alternative that at least one condition differs.
@@ 5.1.3. Wilcoxon Signed Rank Test
For significant Friedman results, post-hoc Wilcoxon signed-rank tests (wilcoxon1992individual, ) assessed pairwise differences (full vs. partial, full vs. no). High skewness (—skewness— ¿ 1 in 11 of 28 comparisons) prompted sign tests to validate Wilcoxon results.
@@ 5.3.2. Model-Specific Heterogeneity
The fact that five of seven models showed no significant effects despite testing on the same code snippets indicates that documentation alone is insufficient for most current LLMs. This contrasts with conventional software engineering wisdom emphasizing documentation importance (boehm1984software, ; sommerville2011software, ), suggesting that LLMs may process code differently than human developers.
@@ 6.1.1. Data Preparation
The evaluation utilized RealClassEval’s Post-Cutoff dataset, ensuring all test cases were from projects created after model training cutoffs. For each target class skeleton, RAG retrieved the top $k=2$ most similar class skeletons from Pre-Cutoff data using cosine similarity on embeddings (obtained from GraphCodeBert (guo2020graphcodebert, )), and then, appended their complete implementations (signature, docstrings, and body) to the prompt. The choice of $k=2$ is motivated by the fact that the length of the prompt can get very long depending on the implementation of the example classes. Since the API service providers of the LLMs charge not only based on the output produced, but also based on the length of input, it was beyond our budget to use $k>2$.
@@ 6.3. Interpretation
The findings reveal a critical interaction between RAG and documentation completeness, answering RQ3: RAG significantly improves code generation quality for unseen tasks, but only when documentation is incomplete. This ”information gap hypothesis” (loewenstein1994psychology, ; golman2018information, ) suggests RAG’s primary value lies in compensating for missing context rather than enhancing already-comprehensive specifications.
@@ 7.1.2. Statistical Analysis
We applied chi-square tests of independence (pearson1900x, ; fisher1970statistical, ) to compare error type distributions across conditions, appropriate for categorical error frequency data. The analysis comprised four complementary comparisons:
Overall Distribution Analysis: Chi-square goodness-of-fit test (conover1999practical, ) against uniform distribution was applied to confirm non-uniform error patterns in order to establish that error types are non-randomly distributed across failures. The larger the chi-squared value, the greater the deviation from the uniform distribution, implying that there is a meaningful pattern worth investigating.
@@ 7.3.1. The AttributeError-TypeError-AssertionError Triad
The absence of SyntaxError (0.00%) confirms that modern LLMs have mastered Python syntax completely. Even though this is opposite to the bug patterns in function-level LLM-generated code reported in a prior study (tambon2025bugs, ), our finding validates LLMs’ use for code generation tasks and shifts attention to semantic correctness—understanding object models, type systems, and logical specifications—as the remaining frontier.
@@ 8.3.4. Methodological Contributions
Real-World Dataset Introduction: RealClassEval provides the first carefully collected real-world class-level benchmark, enabling reproducible evaluation on real code generation tasks. While relatively small, it is 2× larger than ClassEval and represents real industrial software engineering challenges, providing a foundation for future real-world benchmarks.
@@ 9.1.1. Sample Size
The most significant limitation concerns dataset size and the resulting small to negligible effect sizes in the majority of cases. Our study evaluated 200 classes per RealClassEval version, totalling 400 real-world classes compared to ClassEval’s 100 synthetic tasks. While RealClassEval is 2× larger than the established ClassEval benchmark and represents the first carefully collected real-world class-level dataset, it remains relatively small by machine learning evaluation standards. This size limitation appears in consistently small effect sizes across research questions despite statistical significance. Docstring effects and RAG effects showed negligible practical differences, and error distribution comparisons between Pre-Cutoff and Post-Cutoff showed small effects compared to the synthetic-real divide. These small effect sizes reflect three factors: genuine small practical effects, high variability between classes, and limited statistical power to detect small-to-medium effects with 200 samples.
The limited size stems from practical constraints. Generating and testing 18,515 classes across seven LLMs, two dataset versions, three docstring conditions, and two retrieval settings required substantial computational resources. At current API rates of different LLMs’ API service providers, the total generation cost, including trial and error, was approximately \$1000 for API calls alone, plus over 500 CPU hours for test execution. Each additional class requires LLM generation, test suite generation, and test execution, increasing costs directly with size. Expanding to 1,000+ classes per version would require proportional resource scaling, exceeding typical academic research budgets.
@@ 9.1.2. Model Selection and Recency
This study focuses exclusively on class-level code generation, where models must implement complete classes with multiple methods given only skeletons. Findings may not generalize to other granularities: single-function generation may show higher success rates due to a simpler context, while generating entire projects likely shows even larger challenges due to cross-file dependencies. The choice of class-level granularity reflects realistic software engineering tasks and enables controlled evaluation, but complete understanding requires evaluation across multiple granularities. The error pattern insights may not generalize to other object-oriented granularities (tambon2025bugs, ) and require validation.

### v2

@@ 3.1.4. Post-processing and Sampling:
Table 1. An example Python class and corresponding class skeleton from codenerix/django-codenerix (githubGitHubTkaemmingdjangosubdomains).
@@
Start Mine pre-cutoff classes from CodeSearchNet projects Mine post-cutoff classes from repositories created after March 31, 2025 Merge datasets (‘pre’ and ‘post’ flags added for easy tracking) Filter engineered projects Static analysis with Understand™ Extract class skeletons using ast library Remove problematic classes Randomly sample 200 classes per version RealClassEval Dataset
@@ 3.2.1. Selected LLMs
To evaluate the performance of LLMs on class-level code generation, we selected seven state-of-the-art models: Qwen 2.5 Coder (hui2024qwen2), GPT-4.1 (openaiIntroducingGPT41), GPT-5 (openaiIntroducingGPT5), GPT-OSS (agarwal2025gpt), Codestral (mistralCodestralMistral), DeepSeek-V3 (liu2024deepseek), and Llama-4 Maverick (metaLlamaHerd). These models were chosen to represent a diverse range of architectures, training paradigms, and capabilities, ensuring a comprehensive evaluation of both general-purpose and code-specific LLMs, as well as models with varying reasoning abilities. They also vary in their release dates, knowledge cutoffs, parameter sizes, context window lengths, and specialization. Table 2 summarizes these characteristics, highlighting the diversity of our evaluation set.
@@
Qwen 2.5 Coder September 2024 June 2024 32B 128K Code-specific, optimized for code generation and reasoning
@@
Codestral January 2025 (Latest) Unknown 100B 256K Code-specific, specialized for code, lightweight and efficient
@@ 3.2.2. Prompt Design
Unlike ClassEval (du2023classeval), which explores three generation strategies (holistic, incremental, and interactive), we focus exclusively on holistic generation. This decision is motivated by the findings of ClassEval, which indicate that models with longer context windows (such as GPT-3.5 and GPT-4 in their study) perform well in holistic generation. Given that all our selected models have substantial context windows (ranging from 128K to 1M tokens), we opt for holistic generation to leverage their ability to process and generate entire class implementations in a single pass.
@@ 3.3. Test Suite Generation
To assess functional correctness, we use PYNGUIN (lukasczyk2022pynguin) to automatically generate test suites for the classes in our dataset. PYNGUIN (the PYthoN General UnIt test geNerator) is an automated unit test generation framework specifically designed for Python programs. Unlike test generation tools for statically typed languages like Java, PYNGUIN addresses the unique challenges posed by Python’s dynamic typing and lack of explicit type information. The tool takes a Python module as input along with its dependencies and automatically generates unit tests that maximize code coverage using various test generation algorithms.
We specifically employ PYNGUIN’s DynaMOSA (Dynamic Many-Objective Sorting Algorithm) (panichella2017automated) algorithm for test generation. DynaMOSA is a many-objective evolutionary algorithm specifically designed for automated test case generation with dynamic selection of coverage targets. The algorithm treats test case generation as a many-objective optimization problem where each coverage target (e.g., branches, statements) is treated as a separate objective. DynaMOSA dynamically adjusts its focus based on the control dependency hierarchy, prioritizing uncovered targets and progressively exploring the search space. This approach has been shown to achieve higher branch coverage compared to traditional whole-suite or random test generation approaches, making it particularly suitable for evaluating the functional correctness of complex, real-world Python classes.
Given the diverse nature of real-world projects in our dataset, we lack sufficient domain knowledge to manually create meaningful test cases for each class, limiting our ability to measure correctness beyond structural similarity. PYNGUIN with its DynaMOSA algorithm provides a suitable solution as it automatically generates test cases with high branch coverage, enabling us to evaluate the functional correctness of LLM-generated classes without manual intervention. By using automated test generation, we ensure a consistent and reproducible evaluation methodology across all 400 classes in our dataset.
@@ 3.4.1. Performance Analysis
To evaluate the performance of LLMs on class-level code generation, we employ pass rate as our primary metric, defined as the ratio of passed tests to the total number of tests for each generated class. This metric is directly related to the widely used Pass@1 (chen2021evaluating), which measures the probability that a generated solution passes all tests in a single attempt. While Pass@1 is typically used for function-level evaluation, our pass rate extends this concept to class-level correctness, providing a robust measure of functional accuracy across the entire class implementation. For a given class $i$, the pass rate metric is defined as:
@@ 4.1.2. Kruskal-Wallis Test
We used the Kruskal-Wallis test (kruskal1952use) to compare per-class pass rates across the three datasets (ClassEval, Pre-Cutoff, Post-Cutoff) for each LLM. In the current context, the null hypothesis states that the pass rate distributions are identical across datasets, while the alternative hypothesis states that at least one dataset’s distribution differs.
@@
The Kruskal-Wallis test is a non-parametric alternative to one-way ANOVA, suitable for non-normal data and independent samples across three or more groups. The non-normality of pass rates was confirmed by the Shapiro-Wilk test (shapiro1965analysis), justifying the non-parametric approach.
@@ 4.1.3. Mann-Whitney U Tests
For significant Kruskal-Wallis results, we conducted pairwise Mann-Whitney U tests (mann1947test) (ClassEval vs. Pre-Cutoff, ClassEval vs. Post-Cutoff, Pre-Cutoff vs. Post-Cutoff) to identify specific differences. The Mann-Whitney U test is a non-parametric approach suitable for unpaired samples (distinct classes across datasets) and non-normal pass rates. We applied Benjamini-Hochberg (benjamini1995controlling) False Discovery Rate (FDR) correction globally across all 21 tests (3 comparisons per LLM, 7 LLMs) to control for multiple comparisons. The pairwise tests pinpoint which datasets differ, while FDR correction balances Type I error across all tests.
@@ 4.1.4. Effect Sizes and Confidence Intervals
We computed Cliff’s Delta (cliff1993dominance; macbeth2011cliff) to quantify the magnitude of pairwise differences, with thresholds (cohen2013statistical; rahman2024automatic): negligible (¡0.147), small (¡0.33), medium (¡0.474), large ($\geq$0.474). Bootstrap 95% confidence intervals (CIs) were calculated for Cliff’s Delta to quantify uncertainty, using 1,000 iterations. Cliff’s Delta provides an effect size for ordinal data, and CIs ensure robust estimation of differences. Although Cliff’s Delta help us get the effect size of differences, it does not quantify the actual differences. Hence, we also calculate the mean differences between the distributions with bootstrap 95% CIs for an easy interpretation.
@@ 4.1.5. Pooled Analysis
We performed a chi-square analysis (pearson1900x; fisher1922interpretation) on pooled test case counts (passed vs. failed) using 2×2 contingency tables for pairwise dataset comparisons. The chi-square test is appropriate for categorical outcomes. Cramér’s V (cramer1999mathematical) quantified effect sizes (cohen2013statistical): negligible (¡0.1), small (¡0.3), medium (¡0.5), large ($\geq$0.5), with bootstrap CIs for proportion differences. FDR correction was applied globally across all 21 tests. The pooled analysis provides an aggregate view of performance, complementing per-class results and ensuring robustness.
@@ 4.3. Interpretation
The high pass rate of ClassEval compared to RealClassEval was significant, but the reasons were initially unclear, as the large effect sizes (Cliff’s Delta 0.46–0.64) and pass rate differences did not fully explain the gap. To investigate, we conducted a mixed-effects model (laird1982random; carey2001mixed) analysis, including dataset, test count, and LLM-generated code metrics (lines of code, method count, cyclomatic complexity, and class coupling) as predictors, with LLM as a random effect. We chose these $4$ static metrics because they represent both the length of the code as well as the complexity of the code. The model showed significant dataset effects (e.g., $\beta=-1.060$, $p<0.001$) and a negative test count effect ($\beta=-0.014$, $p<0.001$). While lines of code ($\beta=-0.002$, $p=0.006$) and cyclomatic complexity ($\beta=0.043$, $p<0.001$) showed statistical significance, their effects were small relative to the dataset effect, indicating that code complexity metrics explain only a minor portion of the performance gap.
@@ 5.1.2. Friedman Test
Given the continuous and bounded nature of our outcome variable (pass rate $\in[0,1]$) and the paired structure of our data, we employed non-parametric statistical methods. The Friedman test (friedman1937use; friedman1940comparison) compared pass rates across three docstring conditions (full, partial, no) for each model-dataset combination, testing the null hypothesis that pass rate distributions are identical versus the alternative that at least one condition differs.
@@ 5.1.3. Wilcoxon Signed Rank Test
For significant Friedman results, post-hoc Wilcoxon signed-rank tests (wilcoxon1992individual) assessed pairwise differences (full vs. partial, full vs. no). High skewness (—skewness— ¿ 1 in 11 of 28 comparisons) prompted sign tests to validate Wilcoxon results.
@@ 5.3.2. Model-Specific Heterogeneity
The fact that five of seven models showed no significant effects despite testing on the same code snippets indicates that documentation alone is insufficient for most current LLMs. This contrasts with conventional software engineering wisdom emphasizing documentation importance (boehm1984software; sommerville2011software), suggesting that LLMs may process code differently than human developers.
@@ 6.1.1. Data Preparation
The evaluation utilized RealClassEval’s Post-Cutoff dataset, ensuring all test cases were from projects created after model training cutoffs. For each target class skeleton, RAG retrieved the top $k=2$ most similar class skeletons from Pre-Cutoff data using cosine similarity on embeddings (obtained from GraphCodeBert (guo2020graphcodebert)), and then, appended their complete implementations (signature, docstrings, and body) to the prompt. The choice of $k=2$ is motivated by the fact that the length of the prompt can get very long depending on the implementation of the example classes. Since the API service providers of the LLMs charge not only based on the output produced, but also based on the length of input, it was beyond our budget to use $k>2$.
@@ 6.3. Interpretation
The findings reveal a critical interaction between RAG and documentation completeness, answering RQ3: RAG significantly improves code generation quality for unseen tasks, but only when documentation is incomplete. This ”information gap hypothesis” (loewenstein1994psychology; golman2018information) suggests RAG’s primary value lies in compensating for missing context rather than enhancing already-comprehensive specifications.
@@ 7.1.2. Statistical Analysis
We applied chi-square tests of independence (pearson1900x; fisher1970statistical) to compare error type distributions across conditions, appropriate for categorical error frequency data. The analysis comprised four complementary comparisons:
Overall Distribution Analysis: Chi-square goodness-of-fit test (conover1999practical) against uniform distribution was applied to confirm non-uniform error patterns in order to establish that error types are non-randomly distributed across failures. The larger the chi-squared value, the greater the deviation from the uniform distribution, implying that there is a meaningful pattern worth investigating.
@@ 7.3.1. The AttributeError-TypeError-AssertionError Triad
The absence of SyntaxError (0.00%) confirms that modern LLMs have mastered Python syntax completely. Even though this is opposite to the bug patterns in function-level LLM-generated code reported in a prior study (tambon2025bugs), our finding validates LLMs’ use for code generation tasks and shifts attention to semantic correctness—understanding object models, type systems, and logical specifications—as the remaining frontier.
@@ 8.3.4. Methodological Contributions
Real-World Dataset Introduction: RealClassEval provides the first carefully collected real-world class-level benchmark, enabling reproducible evaluation on real code generation tasks. While relatively small, it is 2 times larger than ClassEval and represents real industrial software engineering challenges, providing a foundation for future real-world benchmarks.
@@ 9.1.1. Sample Size
The most significant limitation concerns dataset size and the resulting small to negligible effect sizes in the majority of cases. Our study evaluated 200 classes per RealClassEval version, totalling 400 real-world classes compared to ClassEval’s 100 synthetic tasks. While each split (pre-cutoff and post-cutoff) of RealClassEval is 2 times larger than the established ClassEval benchmark and represents the first carefully collected real-world class-level dataset, it remains relatively small by machine learning evaluation standards. This size limitation appears in consistently small effect sizes across research questions despite statistical significance. Docstring effects and RAG effects showed negligible practical differences, and error distribution comparisons between Pre-Cutoff and Post-Cutoff showed small effects compared to the synthetic-real divide. These small effect sizes reflect three factors: genuine small practical effects, high variability between classes, and limited statistical power to detect small-to-medium effects with 200 samples.
The limited size stems from practical constraints. Generating and testing over 16,000 classes across seven LLMs, two dataset versions, three docstring conditions, and two retrieval settings required substantial computational resources. At current API rates of different LLMs’ API service providers, the total generation cost, including trial and error, was approximately \$1000 for API calls alone, plus over 500 CPU hours for test execution. Each additional class requires LLM generation, test suite generation, and test execution, increasing costs directly with size. Expanding to 1,000+ classes per version would require proportional resource scaling, exceeding typical academic research budgets.
@@ 9.1.2. Model Selection and Recency
This study focuses exclusively on class-level code generation, where models must implement complete classes with multiple methods given only skeletons. Findings may not generalize to other granularities: single-function generation may show higher success rates due to a simpler context, while generating entire projects likely shows even larger challenges due to cross-file dependencies. The choice of class-level granularity reflects realistic software engineering tasks and enables controlled evaluation, but complete understanding requires evaluation across multiple granularities. The error pattern insights may not generalize to other object-oriented granularities (tambon2025bugs) and require validation.